OpenStudy (zzr0ck3r):

Abstract algebra: I have a problem, I need to show that if F is a field then F is an integral domain. easy enough. Let a,b be in F. Assume ab=0 where a is not 0, then since F is a field, we have a^-1ab = b = 0, so every element is a non zero divisor, and so F is an integral domain. my problem RxR is a Field, and (0,1)(1,0) = (0,0), so (1,0) and (0,1) are zero divisors. what am I not seeing?

OpenStudy (anonymous):

Your proof for the first part is correct. However, what you showed afterwards was that \(\Bbb{R}\times\Bbb{R}\) with component wise addition and multiplication has zero divisors and hence \(\Bbb{R}\times\Bbb{R}\) not a field even though \(\Bbb{R}\) itself is a field. Are you sure the question was really claiming that \(\Bbb{R}\times\Bbb{R}\) is a field?

OpenStudy (anonymous):

\(\mathbb{R}^2\) is not a field with componentwise multiplication... if you equip it with \((a,b)\cdot(c,d)=(ac-bd,ad+bc)\) it is however

OpenStudy (anonymous):

^^ that's called the complex numbers \(\mathbb{C}\)

OpenStudy (zzr0ck3r):

word, yeah I thought I had seen a proof that RxR was a field, but of course it cant be (1,0)*(0,1) = (0,0)